# Campana points on wonderful compactifications

### v3

@@ 1. Introduction
More than this however, the original conjectural constant put forward in [PSTVA21] was shown to be incorrect in [Str22] and [Shu22]. We suggest a corrected version of this leading constant (Conjecture 8.3), which we verify to hold in our case. This is given as a (possibility infinite) sum over integrals of local invariants of Brauer group elements with respect to a suitable Tamagawa measure.
@@ 1.1. The main result
We now state our main theorem more precisely. Let $F$ be a number field, ${\mathsf{G}}$ a semi-simple adjoint group over $F$ of rank $\geq 2$, and $X$ the wonderful compactification of ${\mathsf{G}}$ (see [DCP83] for its construction). The boundary divisor $D=X\backslash G$ is a strict normal crossings divisor on $X$. Let $(D_{\alpha})_{\alpha\in\mathcal{A}}$ be its irreducible components defined over $F$. Let $S$ be a finite set of places of $F$ containing all archimedean places such that there is a good integral model $(\mathcal{X},\mathcal{D})$ for $(X,D)$ over the ring of $S$-integers $\mathcal{O}_{F,S}$ of $F$ in the sense of [PSTVA21, Section 3.2]. For each $\alpha\in\mathcal{A},$ let $\epsilon_{\alpha}$ be an element of
@@
in terms of the invariants $a=a((X,D_{\epsilon}),L)),$ $b=b(F,(X,D_{\epsilon}),L)),$ and $c=c(F,S,(\mathcal{X},\mathcal{D}_{\epsilon}),\mathcal{L})$, which are defined in Section 3.3 as well as Conjecture 8.3 when $(X,D_{\epsilon})$ is klt. The main result of this article is the following:
@@
We expect Theorem 1 to be true for the rank $1$ case, but we will not include a proof here, c.f., Remark 6.8.
@@ 1.2. The leading constant
We obtain an explicit expression for the leading constant appearing in Theorem 1. A conjecture for the leading constant for Campana points was put forward in [PSTVA21, Section 3.3]. This conjecture was shown to be incorrect independently by Streeter [Str22] and Shute [Shu22]. The conjectural constant closely mirrored Peyre’s constant [Pey95] from Manin’s conjecture, and the problematic part concerned the factor #$\#\mathrm{H}^{1}(\Gamma,\operatorname{Pic}\bar{U})$ proposed as a replacement for the $\beta$ constant (this part of the constant first appeared in [BT95]).
@@ 2.1. Number fields
Throughout this paper $F$ is a field of characteristic 00. When $F$ is a number field, $\mathcal{O}_{F}$ is the ring of integers of $F$, and $\mathbb{A}_{F}$ is the $F$-algebra of adeles of $F$ endowed with the usual locally compact topology. Let $\Omega_{F}$ be the set of places of $F$, and $\Omega_{F}^{<\infty}$ (resp. $\Omega_{F}^{\infty}$) the set of all nonarchimedean (resp. archimedean) places. For $v\in\Omega_{F}$, $F_{v}$ denotes the completion of $F$ at $v$, and |.$|.|_{v}$ the normalized absolute value on $F_{v}$. For non-archimedean $v$, $\mathcal{O}_{v}$ denotes the ring of integers of $F_{v}$, $\mathfrak{m}_{v}$ the maximal ideal of $\mathcal{O}_{v}$, $k_{v}$ the residue field, $q_{v}$ the order of $k_{v}$, and $\varpi_{v}$ a uniformizer. We write ${\mathbb{A}}_{S}$ for $\prod_{v\notin S}^{\prime}F_{v}$, and ${\mathbb{A}}_{f}$ for ${\mathbb{A}}_{\Omega^{\infty}_{F}}$. For any finite set $S\subset\Omega_{F}$ containing $\Omega_{F}^{\infty}$, let $\mathcal{O}_{F,S}$ be the ring of $S$-integers of $F$, i.e., the set of $a\in F$ such that $|a|_{v}\leq 1$ for all $v\notin S$. Let $\overline{F}$ be an algebraic closure of $F$ and write $\Gamma_{F}$ for the absolute Galois group of $F$.
@@ 2.3. Algebraic groups
There is a perfect pairing $\langle,\rangle:\mathfrak{X}^{*}({\mathsf{S}}_{v})\times\mathfrak{X}_{*}({\mathsf{S}}_{v})\rightarrow\mathbb{Z}$ [Spr98, Lem. .2.11] where $\mathfrak{X}_{*}({\mathsf{S}}_{v})$ denotes the set of cocharacters of ${\mathsf{S}}_{v}$. Since by assumption ${\mathsf{G}}$ is adjoint, the simple roots form a basis of $\mathfrak{X}^{*}({\mathsf{T}})$, and so for each simple root $\theta\in\Delta({\mathsf{G}}_{F_{v}},{\mathsf{S}}_{v})$ we may define a cocharacter $\check{\theta}$ by $\langle\alpha,\check{\theta}\rangle=-\delta_{\alpha\theta},$ where $\delta_{\alpha\theta}$ is 1 if $\alpha=\theta$, and is 0 otherwise.
@@ 2.6. Automorphic characters
By [GMO08, Lemma 4.7(1)] we have #$\#\mathscr{X}^{W_{f}}=[{\mathsf{G}}(\mathbb{A}_{F}):G_{W_{f}}]<\infty.$
@@ 3.1.1. Algebraically closed fields
We review the construction of the wonderful compactification over an algebraically closed field of characteristic 00. The reader can consult [DCP83] for background and details. Let ${\mathsf{G}}$ be a semi-simple linear algebraic group of adjoint type. We first introduce some notation: choose a Borel subgroup ${\mathsf{B}}\subset{\mathsf{G}}$ as well as a maximal torus ${\mathsf{T}}\subset{\mathsf{B}}$. Let $\Phi$ denote the root system of $({\mathsf{G}},{\mathsf{T}})$, $\Phi^{+}\subset\Phi$ the subset of positive roots consisting of roots of ${\mathsf{B}}$, $\Delta=\{\alpha_{1},\dots,\alpha_{l}\}$ the corresponding set of simple roots, $\rho$ the half-sum of positive roots, and $W$ the Weyl group.
The coroot of any $\alpha\in\Phi$ is denoted by $\alpha^{\vee}$; this is a cocharacter of ${\mathsf{T}}$. The coroots form the dual root system $\Phi^{\vee}$. The pairing between characters and cocharacters is denoted by $\langle$,$\rangle$ , ⟩; we have $\langle{\alpha,\alpha^{\vee}\rangle}=2$ for any $\alpha\in\Phi$.
@@ 3.1.2. Non-closed fields
$\operatorname{Eff}^{1}(X)=\bigoplus\mathbb{R}_{\geq 0}[D_{\Gamma_{F}\cdot\beta}]$
where the sum is taken over the $\Gamma$-orbits $\Gamma_{F}\cdot\beta$ in the set $\Delta({\mathsf{G}}^{\prime},{\mathsf{T}}^{\prime})$ of simple roots and $D_{\Gamma\cdot\beta}:=\sum_{\beta^{\prime}\in\Gamma\cdot\beta}D_{\beta^{\prime}}$. As a consequence the set $\mathcal{A}$ in Section 1.1 can be identified with the set of Galois orbits on $\Delta({\mathsf{G}}^{\prime},{\mathsf{T}}^{\prime})$. The anticanonical class is given by
$-K_{X}=\sum_{\alpha\in\mathcal{A}}(\kappa_{\alpha}+1)D_{\alpha},$ (3.1)
@@ 3.3.1. Log Manin’s conjecture for Fano orbifolds
Let $(X,D_{\epsilon}=\sum_{\alpha\in\mathcal{A}}\epsilon_{\alpha}D_{\alpha})$ be a Campana orbifold defined over a number field $F$. We say it is a Fano orbifold if $-(K_{X}+D_{\epsilon})$ is ample.
@@
[…]
containing $\text{containing }a((X,D_{\epsilon}),L)[L]+[K_{X}+D_{\epsilon}].$ (3.4)
@@ 4. Height functions
Let $(X,L)$ be a variety over a number field $F$ together with a line bundle $L$ on $X$. For a place $v\in\Omega_{F}$, a $v$-adic metric on $L$ is a map which assigns to every point $x_{v}\in X(F_{v})$ a function ||.$||.||_{v}:L(x_{v})\rightarrow\mathbb{R}_{\geq 0}$ on the fiber of $L$ at $x_{v}$ such that
@@
    For any open subset $U\subset X$ and any local section $s\in\Gamma(U,L)$, the function on $U(F_{v})$ given by $x_{v}\mapsto||s(x_{v})||_{v}$ is continuous in the $v$-adic topology.
We say a $v$-adic metric on $L$ is smooth if for any open subset $U\subset X$ and any nowhere vanishing section $s\in\Gamma(U,L)$, the function on $U(F_{v})$ given by $x_{v}\mapsto||s(x_{v})||_{v}$ is $C^{\infty}$ if $v$ is archimedean and is locally constant if $v$ is non-archimedean.
@@
Let $(X,L)$ be a projective variety over a number field $F$ together with a line bundle $L$ on $X$. A (smooth) adelic metric on $L$ is a collection ||.||={||.$||.||=\{||.||_{v}\}_{v\in\Omega_{F}}$ of (smooth) $v$-adic metrics for each place $v\in\Omega_{F}$, such that all but finitely many of the ||.$||.||_{v}$ are defined by a single projective model of $(X,L)$ over $\mathcal{O}_{F,S}$ for some $S$ a finite set of $\Omega_{F}$ including all archimedean places.
@@ 4.2. Height functions on the wonderful compactification
$\mathsf{H}_{v}:\textrm{Pic}(X)_{\mathbb{C}}\times{\mathsf{G}}(F_{v})\rightarrow\mathbb{C}^{\times},\ \left(\sum_{\alpha\in\mathcal{A}}s_{\alpha}D_{\alpha},x\right)\mapsto\prod_{\alpha\in\mathcal{A}}||\mathsf{f}_{\alpha}(x)||_{v}^{-s_{\alpha}}.$
This pairing varies linearly on the factor Pic$(X)_{\mathbb{C}}$ and continuously on the factor ${\mathsf{G}}(F_{v})$. We define the global height pairing ${\mathsf{H}}$ as the product of the local height pairings
@@
[…]
@@ 4.3. Height zeta function
$+\sum_{\chi\in\mathfrak{X}}\sum_{P}n(A_{P})^{-1}\int_{\Pi(M_{P})}(\sum_{\phi\in B_{P}(\pi)_{\chi}}E(e,\phi)\int_{G(\mathbb{A}_{F})}\delta_{\epsilon}(g)\mathsf{H}(\mathbb{s},g)^{-1}\overline{E(g,\phi)}\ \mathrm{d}g)\ \mathrm{d}\pi_{p}.$
@@
[…]
with ${\boldsymbol{\Delta}},\|\,.\,\|$ as in [STBT07, Section 3.2], converge. These integrals still converge if we replace $\mathsf{H}(\mathbb{s},g)$ by $\mathsf{H}(\mathbb{s},g)\delta_{\epsilon}(g)$. We now apply [STBT07, Lemma 3.1] to get the result. ∎
@@ 5. Height integrals I: one-dimensional representations
$\int_{{\mathsf{G}}(\mathbb{A}_{F})}\delta_{\epsilon}(g)\mathsf{H}(\mathbb{s},g)^{-1}\chi(g)\,\mathrm{d}g=\prod_{v\in\Omega_{F}}\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g)^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}$
@@
$\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g)^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}=0.$
@@
The assumption $\operatorname{Re}\mathbf{s}\gg 0$ guarantees the absolute convergence of the integral. Pick $\kappa_{v}\in{\mathsf{K}}_{v}$ such that $\chi_{v}(\kappa_{v})\neq 1$. Then making a change of variable $g\mapsto g\kappa$ and using the ${\mathsf{K}}_{v}$-invariance of $\mathsf{H}_{v}$ and $\delta_{\epsilon,v}$, gives
$\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g)^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}=\chi_{v}(\kappa_{v})\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g)^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}.$
@@
In particular we need only consider automorphic characters $\chi=\otimes^{\prime}_{v}\chi_{v}\in\mathscr{X}^{{\mathsf{K}}_{0}}$ where $\mathscr{X}^{{\mathsf{K}}_{0}}$ is the group of automorphic characters which are invariant under the action of ${\mathsf{K}}_{0}$.

### v4

@@ 1. Introduction
More than this however, the original conjectural constant put forward in [PSTVA21] was shown to be incorrect in [Str22] and [Shu22]. We suggest a corrected version of this leading constant (Conjecture 8.3), which we verify to hold in our case. This is given as a (possibly infinite) sum over integrals of local invariants of Brauer group elements with respect to a suitable Tamagawa measure.
@@ 1.1. The main result
We now state our main theorem more precisely. Let $F$ be a number field, $\mathsf{G}$} a semi-simple adjoint group over $F$ of rank $\geq 2$, and $X$ the wonderful compactification of $\mathsf{G}$} (see [DCP83] for its construction). The boundary divisor $D=X\backslash{\mathsf{G}}$ is a strict normal crossings divisor on $X$. Let $(D_{\alpha})_{\alpha\in\mathcal{A}}$ be its irreducible components defined over $F$. Let $S$ be a finite set of places of $F$ containing all archimedean places such that there is a good integral model $(\mathcal{X},\mathcal{D})$ for $(X,D)$ over the ring of $S$-integers $\mathcal{O}_{F,S}$ of $F$ in the sense of [PSTVA21, Section 3.2]. For each $\alpha\in\mathcal{A},$ let $\epsilon_{\alpha}$ be an element of
@@
in terms of the invariants $a=a((X,D_{\epsilon}),L),$ $b=b(F,(X,D_{\epsilon}),L),$ and $c=c(F,S,(\mathcal{X},\mathcal{D}_{\epsilon}),\mathcal{L})$, which are defined in Section 3.3 as well as Conjecture 8.3 when $(X,D_{\epsilon})$ is klt. The main result of this article is the following:
@@
We expect Theorem 1 to be true for the rank 11 case, but we will not include a proof here, c.f., Remark 6.8.
@@ 1.2. The leading constant
We obtain an explicit expression for the leading constant appearing in Theorem 1. A conjecture for the leading constant for Campana points was put forward in [PSTVA21, Section 3.3]. This conjecture was shown to be incorrect independently by Streeter [Str22] and Shute [Shu22]. The conjectural constant closely mirrored Peyre’s constant [Pey95] from Manin’s conjecture, and the problematic part concerned the factor $\mathrm{H}^{1}(\Gamma_{F},\operatorname{Pic}\bar{U})$ proposed as a replacement for the $\beta$ constant (this part of the constant first appeared in [BT95]).
@@ 2.1. Number fields
Throughout this paper $F$ is a field of characteristic 0. When $F$ is a number field, $\mathcal{O}_{F}$ is the ring of integers of $F$, and $\mathbb{A}_{F}$ is the $F$-algebra of adeles of $F$ endowed with the usual locally compact topology. Let $\Omega_{F}$ be the set of places of $F$, and $\Omega_{F}^{<\infty}$ (resp. $\Omega_{F}^{\infty}$) the set of all nonarchimedean (resp. archimedean) places. For $v\in\Omega_{F}$, $F_{v}$ denotes the completion of $F$ at $v$, and $|.|_{v}$ the normalized absolute value on $F_{v}$. (See [PSTVA21, Section 2.1] for a way to normalize absolute values.) For non-archimedean $v$, $\mathcal{O}_{v}$ denotes the ring of integers of $F_{v}$, $\mathfrak{m}_{v}$ the maximal ideal of $\mathcal{O}_{v}$, $k_{v}$ the residue field, $q_{v}$ the order of $k_{v}$, and $\varpi_{v}$ a uniformizer. We write ${\mathbb{A}}_{S}$ for $\prod_{v\notin S}^{\prime}F_{v}$, and ${\mathbb{A}}_{f}$ for ${\mathbb{A}}_{\Omega^{\infty}_{F}}$. For any finite set $S\subset\Omega_{F}$ containing $\Omega_{F}^{\infty}$, let $\mathcal{O}_{F,S}$ be the ring of $S$-integers of $F$, i.e., the set of $a\in F$ such that $|a|_{v}\leq 1$ for all $v\notin S$. Let $\overline{F}$ be an algebraic closure of $F$ and write $\Gamma_{F}$ for the absolute Galois group of $F$.
@@ 2.3. Algebraic groups
There is a perfect pairing $\langle,\rangle:\mathfrak{X}^{*}({\mathsf{S}}_{v})\times\mathfrak{X}_{*}({\mathsf{S}}_{v})\rightarrow\mathbb{Z}$ [Spr98, Lem. 2.11] where $\mathfrak{X}_{*}({\mathsf{S}}_{v})$ denotes the set of cocharacters of ${\mathsf{S}}_{v}$. Since by assumption $\mathsf{G}$} is adjoint, the simple roots form a basis of $\mathfrak{X}^{*}({\mathsf{T}})$, and so for each simple root $\theta\in\Delta({\mathsf{G}}_{F_{v}},{\mathsf{S}}_{v})$ we may define a cocharacter $\check{\theta}$ by $\langle\alpha,\check{\theta}\rangle=-\delta_{\alpha\theta},$ where $\delta_{\alpha\theta}$ is 1 if $\alpha=\theta$, and is 0 otherwise.
@@ 2.6. Automorphic characters
By [GMO08, Lemma 4.7(1)] we have $\mathscr{X}^{W_{f}}=[{\mathsf{G}}(\mathbb{A}_{F}):G_{W_{f}}]<\infty.$
@@ 3.1.1. Algebraically closed fields
We review the construction of the wonderful compactification over an algebraically closed field of characteristic 0. The reader can consult [DCP83] for background and details. Let $\mathsf{G}$} be a semi-simple linear algebraic group of adjoint type. We first introduce some notation: choose a Borel subgroup ${\mathsf{B}}\subset{\mathsf{G}}$ as well as a maximal torus ${\mathsf{T}}\subset{\mathsf{B}}$. Let $\Phi$ denote the root system of $({\mathsf{G}},{\mathsf{T}})$, $\Phi^{+}\subset\Phi$ the subset of positive roots consisting of roots of $\mathsf{B}$}, $\Delta=\{\alpha_{1},\dots,\alpha_{l}\}$ the corresponding set of simple roots, $\rho$ the half-sum of positive roots, and $W$ the Weyl group.
The coroot of any $\alpha\in\Phi$ is denoted by $\alpha^{\vee}$; this is a cocharacter of $\mathsf{T}$}. The coroots form the dual root system $\Phi^{\vee}$. The pairing between characters and cocharacters is denoted by $\langle,\rangle$; we have $\langle{\alpha,\alpha^{\vee}\rangle}=2$ for any $\alpha\in\Phi$.
@@ 3.1.2. Non-closed fields
$\operatorname{Eff}^{1}(X)=\bigoplus\mathbb{R}_{\geq 0}[D_{\Gamma\cdot\beta}]$
where the sum is taken over the $\Gamma$-orbits $\Gamma\cdot\beta$ in the set $\Delta({\mathsf{G}}^{\prime},{\mathsf{T}}^{\prime})$ of simple roots and $D_{\Gamma\cdot\beta}:=\sum_{\beta^{\prime}\in\Gamma\cdot\beta}D_{\beta^{\prime}}$. As a consequence the set $\mathcal{A}$ in Section 1.1 can be identified with the set of Galois orbits on $\Delta({\mathsf{G}}^{\prime},{\mathsf{T}}^{\prime})$. The anticanonical class is given by
$K_{X}=\sum_{\alpha\in\mathcal{A}}(\kappa_{\alpha}+1)D_{\alpha},$ (3.1)
@@ 3.3.1. Log Manin’s conjecture for Fano orbifolds
Let $(X,D_{\epsilon}=\sum_{\alpha\in\mathcal{A}}\epsilon_{\alpha}D_{\alpha})$ be a Campana orbifold defined over a number field $F$. We say it is a Fano orbifold if $(K_{X}+D_{\epsilon})$ is ample.
@@
$b(F,(X,D_{\epsilon}),L)=$ $\text{ the codimension of the minimal face of }\operatorname{Eff}^{1}(X)$ (3.3)
$\text{containing }a((X,D_{\epsilon}),L)[L]+[K_{X}+D_{\epsilon}].$
@@ 4. Height functions
Let $(X,L)$ be a variety over a number field $F$ together with a line bundle $L$ on $X$. For a place $v\in\Omega_{F}$, a $v$-adic metric on $L$ is a map which assigns to every point $x_{v}\in X(F_{v})$ a function $||.||_{v}:L(x_{v})\rightarrow\mathbb{R}_{\geq 0}$ on the fiber of $L$ at $x_{v}$ such that
@@
[…]
@@
Let $(X,L)$ be a projective variety over a number field $F$ together with a line bundle $L$ on $X$. A (smooth) adelic metric on $L$ is a collection $||.||=\{||.||_{v}\}_{v\in\Omega_{F}}$ of (smooth) $v$-adic metrics for each place $v\in\Omega_{F}$, such that all but finitely many of the $||.||_{v}$ are defined by a single projective model of $(X,L)$ over $\mathcal{O}_{F,S}$ for some $S$ a finite set of $\Omega_{F}$ including all archimedean places.
@@ 4.2. Height functions on the wonderful compactification
[…]
@@
From now on we will use notations introduced in Section 2.3. For all but finitely many places $v$ of $F$, the group ${\mathsf{G}}_{F_{v}}$ is quasi-split [PR94, Thm. 6.7]. Let $v$ be a place of $F$ for which ${\mathsf{G}}_{F_{v}}$ is quasi-split. Then each tuple $\mathbb{s}=(s_{\alpha})_{\alpha\in\mathcal{A}}$ gives rise to a tuple $\mathbb{s}=(s^{v}_{\theta})_{\theta\in\Delta({\mathsf{G}}_{F_{v}},{\mathsf{S}}_{v})}$. Indeed, when $\theta$ is above $\alpha$, we define as $s^{v}_{\theta}=s_{\alpha}$. When there is no danger of confusion we simply write $\mathbf{s}=(s_{\theta})_{\theta\in\Delta({\mathsf{G}}_{F_{v}},{\mathsf{S}}_{v})}$. We use these complex variables to perform our local computations without comment.
@@ 4.3. Height zeta function
$+\sum_{\chi\in\mathfrak{X}}\sum_{P}n(A_{P})^{-1}\int_{\Pi(M_{P})}\left(\sum_{\phi\in B_{P}(\pi)_{\chi}}E(e,\phi)\int_{G(\mathbb{A}_{F})}\delta_{\epsilon}(g)\mathsf{H}(\mathbb{s},g)^{-1}\overline{E(g,\phi)}\ \mathrm{d}g\right)\ \mathrm{d}\pi_{p}.$
@@
Note that the notation in the above formula is explained in [LTBT20, Section 4.4], and we do not repeat those in this paper.
@@
$\int_{{\mathsf{G}}(\mathbb{A}_{F})}|\mathsf{H}(\mathbb{s},g)|\|g\|^{N}\,dg,\quad\int_{{\mathsf{G}}(\mathbb{A}_{F})}|{\boldsymbol{\Delta}}^{n}\mathsf{H}(\mathbb{s},g)|\|g\|^{N}\,dg,$
with $\boldsymbol{\Delta}$$\|\,$$\|$ as in [STBT07, Section 3.2], converge. These integrals still converge if we replace $\mathsf{H}(\mathbb{s},g)$ by $\mathsf{H}(\mathbb{s},g)\delta_{\epsilon}(g)$. We now apply [STBT07, Lemma 3.1] to get the result. ∎
@@ 5. Height integrals I: one-dimensional representations
$\int_{{\mathsf{G}}(\mathbb{A}_{F})}\delta_{\epsilon}(g)\mathsf{H}(\mathbb{s},g)^{-1}\chi(g)\,\mathrm{d}g=\prod_{v\in\Omega_{F}}\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g_{v})^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}$
@@
$\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g_{v})^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}=0.$
@@
The assumption $\operatorname{Re}\mathbf{s}\gg 0$ guarantees the absolute convergence of the integral. Pick $\kappa_{v}\in{\mathsf{K}}_{v}$ such that $\chi_{v}(\kappa_{v})\neq 1$. Then making a change of variable $g_{v}\mapsto g_{v}\kappa_{v}$ and using the ${\mathsf{K}}_{v}$-invariance of $\mathsf{H}_{v}$ and $\delta_{\epsilon,v}$, gives
$\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g_{v})^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}=\chi_{v}(\kappa_{v})\int_{{\mathsf{G}}(F_{v})}\mathsf{H}_{v}(\mathbb{s},g_{v})^{-1}\chi_{v}(g_{v})\delta_{\epsilon,v}(g_{v})\,\mathrm{d}g_{v}.$
@@
In particular we need to only consider automorphic characters $\chi=\otimes^{\prime}_{v}\chi_{v}\in\mathscr{X}^{{\mathsf{K}}_{0}}$ where $\mathscr{X}^{{\mathsf{K}}_{0}}$ is the group of automorphic characters which are invariant under the action of ${\mathsf{K}}_{0}$ and $\otimes^{\prime}_{v}$ means the restricted tensor product.